OpenStudy (anonymous):

what are the foci of the graph 4X^2-9Y=36? ....

OpenStudy (dumbcow):

solve for y y = 4/9 x^2 -4 vertex: (0,-4) foci is -4 + 1/(4*4/9) foci: (0,-55/16)

OpenStudy (anonymous):

possible answers are ((plus minus)6,0), ((plusminus) 3.6, )) , (0, (plus minus)6) , or (0, (plusminus 3.6))

OpenStudy (dumbcow):

hmm..this must be a hyperbola then you mistyped the question 4x^2 -9y^2 = 36 x^2/9 - y^2/4 = 1 center: (0,0) a^2=9 b^2=4 c^2 = 9+4=13 c = sqrt(13) = 3.6 foci: (0, +-3.6)